OpenStudy (anonymous):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1/3 π (x - 7)2 (x - 3) What does the factor 1/3 (x - 7) represent?

OpenStudy (anonymous):

r = 4 - h r = 4 - (x - 3) r = 4 - x + 3 7 - x = r so there is where your x - 7 is.....

OpenStudy (anonymous):

x-7 is your radius

OpenStudy (anonymous):

\[V_{cone}=\frac{1}{3} \pi r^2 \cdot h\] where r = radius and h = height

OpenStudy (anonymous):

so can you think of why 1/3

OpenStudy (anonymous):

think of how a cone looks like

OpenStudy (anonymous):

 now try to mathmatically explain why a cone's volume is 1/3 of what a cyllinders' is

OpenStudy (anonymous):

So 1/3 (x - 7) represents the total surface area of the tank? There are my options: 1.) The total surface area of the tank 2.) One third of the radius of the tank 3.) One third of the height of the tank 4.) The total area of the base of the tank

OpenStudy (anonymous):

i have class so maybe someone like @ParthKohli can help you, but if you look up like "Definition of a Cone or proof of a cone you can find it

OpenStudy (anonymous):

check this site out... http://nrich.maths.org/1408